In 1895, the first putting green championship was held. The winners prize money was $200. In 2014, the winners check was 1,400,000. What was the percentage increase per year in the winners check over this period? If the winners prize increases at the same rate, what will it be in 2035?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

1,400,000=200*(1+r/100)^119

(1,400,000/200)^(1/119)=(1+r/100)

(1+r/100)=1.0772

r=1.0772-1

=7.72%(Approx)=percentage increase per year.

Hence Value in 2035=1,400,000*(1.0772)^21

=$6678435.70(Approx).